A binomial probability experiment is conducted with the given parameters. Compute the probability of x successes in the n independent trials of the experiment. The probability of x less than or equals 3 is ____.

n=9​, p =0.9​, x less than or equals 3.
Answer #1

solution

Given that ,

p = 0.9

1 - p = 1 - 0.9 = 0.1

n = 9

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X < 3 ) = P(X=0) +P(X=1)+P(X=2)+P(X=3)

= (9 C 0) * 0.90 * (0.1)9+(9 C 1) * 0.91 * (0.1)8+(9 C 2) * 0.92 * (0.1)7+(9 C 3) * 0.93 * (0.1)6

probability =0.00006